OpenStudy (anonymous):

Multiple choice! Answer please?? One of the factors of the polynomial x³ + 5x^2 + 6x is (x^2 + 3x). What is the other factor? A.x-2 B.x+2 C.X+1 D. x-1

OpenStudy (anonymous):

@ehuman

OpenStudy (anonymous):

@LovelyAnna

OpenStudy (amoodarya):

divide it by (x^2 + 3x) to find the answer

OpenStudy (amoodarya):

x³ + 5x^2 + 6x another way is x( x^2 + 5x + 6)= x(x+2)(x+3) now divide by x(x+3)

OpenStudy (amoodarya):

x³ + 5x^2 + 6x x(x+2)(x+3) -------------- = ------------- can you go on ? x^2+3x x(x+3)

OpenStudy (anonymous):

b?

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

thanks